. For each of the following demand functions, determine if the good is normal or inferior. If it is normal, is it a necessary, or luxury good? x1 = 1000 + p2 √ m − p1 √ m
Answer #1

x1 =1000 + p2\sqrt{}m - p1 \sqrt{} m

x1 =1000+ \sqrt{} m (p2- p1)

Take the first order differentiation of x1 w.r.t m , we get:

= 1/2 (p2- p1) m-1/2

Income elasticity of demand = 1/2 (p2- p1) m-1/2 (m/x1> 0.

Because income elasticity is greater than 0 or positive .This implies that the good is a normal good.

The income elasticity is less than 1 , therefore ,it is a necessity good.

If the income elasticity is more than 1 ,then it would be a luxury good.


answered by: ANURANJAN SARSAM